OpenStudy (anonymous):

7 -3 = 7 + -3.....what property is that?

OpenStudy (anonymous):

you change the subtraction operation to addition............usually thats inverse property.........but it might be closure property.....how can you distinguish

OpenStudy (anonymous):

please help

Directrix (directrix):

Let's take a look at some properties and see what we can figure out. Open the attachment.

OpenStudy (anonymous):

ok...thx

Directrix (directrix):

I did not see anything there. Is the operation of subtraction defined in your book?

OpenStudy (anonymous):

but i dont see the property of my question

OpenStudy (anonymous):

no....not really they said to put a propery

Directrix (directrix):

7 -3 = 7 + -3 --> I don't know a named property for that. 7 minus 3 is, by definition of subtration, 7 + (-3). [Change the sign of the second number and add it to the first: 7 -3 = 7 - (+3) = 7 - 3. Sorry to be of no help.